OpenStudy (chris215):

help! the integral of 4^(3x) dx

OpenStudy (chris215):

I keep getting 4^3x/ 3ln(4) +C

OpenStudy (chris215):

but that's not an answer choice

OpenStudy (idku):

\(\color{black}{\displaystyle y=a^x}\) \(\color{black}{\displaystyle \ln y=\ln a^x}\) \(\color{black}{\displaystyle \ln y=x \ln a}\) \(\color{black}{\displaystyle y'/y= \ln a}\) \(\color{black}{\displaystyle y'= y\ln a}\) \(\color{black}{\displaystyle y'= a^x\ln a}\) Since the derivative of \(a^x\) is \(a^x\ln a\), therefore, the antiderivative of \(a^x\) is \((a^x/\ln a)\color{gray}{+C}\). So, you may conclude, \(\color{black}{\displaystyle \int a^x~dx=\frac{a^x}{\ln a} +C:\quad \quad \left\{\left.a\right|a>0,~a\ne e\right\}}\)

OpenStudy (idku):

Well, a can also be equal to e, actually.... e^x/ln(e) = e^x/1 = e^x, which just repeats that integral of e^x is just e^x

OpenStudy (freckles):

try using properties of log @chris215

OpenStudy (freckles):

maybe like the power rule

OpenStudy (freckles):

\[\ln(x^r)=? \text{ thing power rule for log}\]

OpenStudy (idku):

\(\color{black}{\displaystyle \int b^{cx}~dx=\int (b^c)^{x}~dx=\frac{(b^c)^x}{\ln (b^c)} +C=\frac{b^{cx}}{c\ln (b)} +C}\)

OpenStudy (idku):

this is another derivation from my previous derivation, using a property \(a^{bc}=(a^b)^c\)

OpenStudy (chris215):

I got 4^(3x)/(ln4)+C

OpenStudy (freckles):

ln(x^r)=r ln(x)

OpenStudy (freckles):

so 3 ln(4)=?

OpenStudy (chris215):

6ln2?

OpenStudy (chris215):

so it would be 4^(3x)/6ln2 +c ?

OpenStudy (freckles):

you could write 3 ln(4) as 6 ln(2)

OpenStudy (idku):

6•ln(2) = 4•ln(3) Yes

OpenStudy (freckles):

is that one of your choices ?

OpenStudy (idku):

oh the other way, 3 ln(4)

OpenStudy (idku):

yes, you are right.

OpenStudy (chris215):

yes it is thanks so much !

OpenStudy (freckles):

lol I didn't even notice with my dislexic-ness that what you wrote was different than what I wrote @idku

OpenStudy (freckles):

but after much concentration i finally seen it

OpenStudy (idku):

:)